# Fatigue-PINN: Physics-Informed Fatigue-Driven Motion Modulation and Synthesis

### v1

@@ III-A Dataset & Data Processing
Fatigue-PINN leverages a PINN version of 3CC, namely 3CC-$\lambda$, for modeling the fatigue state of specific muscle groups/joint actuator torques (w.r.t. the animated motion) during open human movements. As shown in Figure 1, our PINN 3CC-$\lambda$ is given as input the current active motor units $M_{A}$, which in our case are the joint torques of each DoF, $\hat{\mathbf{\tau}}_{t}^{N}$, produced by the ID model, and outputs the percentage of motor units in the complementary states, $M_{F}$ and $M_{R}$ in $MVC\%$. In contrast to previous work [44], we compute the $RC$ of each joint, based on the $M_{F}$ state multiplied by a $\lambda$ factor (Equation (7)) that empirically describes the fatigue profile of each muscle/joint to better align joint fatigue in different movements and allow for joint-specific fatigue configurations. Subsequently, we apply $RC$ as a time-varying factor to simulate the gradual decaying of the maximum exerted joint torques to directly generate the fatigued ones, $\acute{\hat{\mathbf{\tau}}}_{t}^{N}$, as in [23, 44] (Equation (8)):
@@
For instance, if $M_{F}=20\%$, at elbow joint at frame t, and $\lambda=0.6$ then from Equation (7) we derive that the maximum exerted elbow torques peak at $RC(t)=88\%$ (instead of $RC(t)=80\%$), meaning that $\lambda$ factor, contributes in smoothly injecting joint-specific fatigue and producing more physically-plausible motions without intermittent (”step-like”) movements.
As depicted in Figure 1, the 3CC-$\lambda$ model consists of five fully connected layers, being fed with the current time frame $t$ and current torque of each DoF, $\hat{\mathbf{\tau}}_{t}^{N}$ (i.e. $M_{A}$), and produces $M_{F}$ and $M_{R}$ of the same frame. Its loss, $L$ as formulated in Equations (9), (10), and (11), is comprised of two terms (in accordance with Figure 1) i) the Mean Squared Error (MSE) loss of the neural network, $L_{NN}$, and ii) the physics-based loss, $L_{PB}$, based on Equations (3) and (1) that describe the transition between the active and fatigued state and the transition between the fatigued and the resting state, respectively:
@@ III-C Inverse/Forward Dynamics Model
The forward and backward sequence outputs are computed using Equations (12) and (13), which are the standard LSTM updating equations [66, 68, 67]. In particular, the output $\overrightarrow{h}_{t}$ of the forward sequence is computed iteratively in a positive time direction spanning from $t=T$ to time $t=1$. On the contrary, the backward loop is ”looking” at the past, thus, its output $\overleftarrow{h}_{t}$ is calculated in a negative time direction from $t=1$ to $t=T$ (see Figure 3):
@@
where $x_{t}$, is the input data $x$ at time frame $t$, $U$ and $W$ are the weight matrices and $b$ is the bias vector. The input of the ID model, is defined as $x_{t}=\mathbf{q}_{t}^{N}$, i.e. a vector of $N$ joint angles at frame $t$, while $x_{t}=\hat{\mathbf{\tau}}_{t}^{N}$ is the input for the FD model, i.e. a vector of $N$ fatigued torques at time $t$. The values of the weight matrices and the bias vector are updated (”learned”) during training.
Furthermore, the BiLSTM layer produces an output vector at each time frame $t$, $y_{t}$, which is computed using the calculated feedback sequences’ outputs, $\overrightarrow{h}_{t}$ and $\overleftarrow{h}_{t}$, as indicated by Equation (14) [67].
@@
where in our case $g$ is a concatenating function used to combine the outputs of the two feedback sequences [68]. Given that $y_{t}$, is the output of the last BiLSTM layer in each of ID/FD architecture, it is then passed to a fully connected layer (as illustrated in Figures 1 and 3). In our joint-specific framework, the fully connected layer in the ID model computes the current torque for one joint denoted as $\hat{\mathbf{\tau}}_{t}$, while the fully connected layer of the FD architecture produces a single fatigued joint angle, $\acute{\hat{\mathbf{q}}}_{t}$, at frame $t$.
As shown in Figures 1 and 3, the ID and FD models utilize a symmetric BiLSTM architecture consisting of five BiLSTM layers each one of them containing 128 units with linear activation on the first BiLSTM layer and ReLU activations on the rest hidden layers. Additionally, a Fully Connected layer with linear activation is included as the output of the ID/FD models. Similarly to the training of 3CC-$\lambda$, both ID and FD BiLSTM networks were trained in Python Keras, utilizing the Adam optimizer with a learning rate of $0.001$. The learning process was conducted with a batch size of 32 and a MSE loss function, over 1000 epochs, with early stopping enabled to prevent the model from over-fitting.
@@ IV Results
Figure 4: The impact of different levels of fatigue on punching (Frame 290, i.e. the moment of hit), throwing (Frame 240, at throw), and waving (Frame 100) motions. The levels of fatigue are defined w.r.t Residual Capacity as arises from Equation (8). For instance, $30\%$ fatigue $\rightarrow RC(t)=70\%$, etc.
@@ IV-A Fatigue Effects in Animation
The proposed Fatigue-PINN framework was trained and tested upon open-type movements (as defined in this work), i.e. waving, punching, throwing, to assess its capability of effectively producing fatigued motion in various action classes (see Figure 4). We opted to apply fatigued torques on the upper body (shoulders, elbows, wrists, and lumbar), whose motion is crucial for perceiving the effects of fatigue in these particular open motions. Our results are simulated in Unity Game Engine Environment [69], utilizing the Anime Girl Character [^3] Available online at: https://assetstore.unity.com/packages/3d/characters/humanoids/casual-1-anime-girl-characters-185076#publisher avatar from Unity Asset Store.
@@
In both motion frames depicted in Figure 5, the character on the left exhibits fatigued movement ($RC(t)=\sim 55\%$) in all joints of both arms, while the right is the non-fatigued one. In punching motion (Figure 5), the hand positioning is different than in the non-fatigued motion, rendering the fatigued character ”unguarded” to hits (see frame $100$ in Figure 5). Moreover, in frame $290$ as shown in the right section of Figure 5, another discrepancy in motion due to fatigue is apparent, with both arms performing a wider movement while hitting, thus, missing the target. These two effects on the kinesiology of the 3D character, guard drop, and non-accurate punches, are behaviors observed alongside the increase of perceived fatigue in boxing athletes [18, 19], while guard drops pose as a tactical strategy to mitigate fatigue by providing a small period of rest [18]. The latter, known as a pacing strategy, involves the adaptation of exercise intensity (decrease/increase effort accordingly) to minimize the impact of fatigue and thereby attain the desired outcome [17] (e.g. determine the winner in a boxing bout). Moreover, the ”unguarded” stance is a result of increased shoulder abduction, therefore, decreased shoulder adduction as seen in Section A of Figure 9, and increased elbow extension (i.e. less flexed elbow as in Figure 9). Notably, while shoulder adduction generally decreases, it increases at the moment of hitting impact (around frame $250$), resulting in a corresponding decrease in shoulder abduction. These findings align with those in [3]. Lumbar flexion and rotation are reduced, with the character inclining to the left during striking, while the upper body exhibits greater rotation to the right. These discrepancies in joint angles are reflected in the respective torques, which are expected to be reduced in presence of fatigue by the 3CC-$\lambda$ model.
@@
Similar results are observed in throwing (see Figure 6), with the right hand, the one performing the motion, being directed in a lower position than the one intended. The externally perceived impact of fatigue in this movement can be associated with the one in pointing tasks [70, 2], where one hand extends towards a specific position while the other remains relatively inactive. However, throwing is inherently more physically demanding, thereby we intuitively expect it to lead to a more pronounced fatigue effect. Regarding the changes in position as depicted in the right section of Figure 6 (frame $240$), the shoulder joint positions of both arms are moving towards the non-throwing side, while the right shoulder is positioned higher at throw, which is in accordance with the findings in [70], and may provide compensation mechanisms for the increased shoulder elevation and elbow extension [70, 71]. Moreover, following fatigue, the shoulder and elbow joints of the right arm (the one performing the throwing/pointing task) adopt a more posterior position, indicating a posture that reduces shoulder torques [70] (Figure 7), consistent with the predictions of the 3CC-$\lambda$ model.
Similar to the findings in [70, 2], the fatigued character in the left of Figure 7, exhibits reduced shoulder flexion of the right hand at the moment of throwing (frame $240$). Akin to the results in [70], upper-body inclination towards the non-reaching arm’s side due to decreased lumbar bending angle (i.e. more flexion in the lateral plane towards the left side) is also evident at throw (Frame $240$ in Figure 6). Moreover, in accordance to [2], greater lumbar rotation (i.e. increased lumbar left rotation angle) throughout throwing (frames $210-300$) is apparent in fatigued movement as illustrated in Figures 6 and 9 (Section B). The above kinematic effects are attributed to elbow and/or shoulder fatigue, as observed in [70, 1, 2].
Furthermore, the right-hand elbow is less flexed (reduced elbow flexion angle as seen in Figures 6 and 9), which is considered a mean of compensation for both elbow and trunk fatigue as mentioned in [2], whereas elbow fatigue also influences shoulder abduction/adduction, leading to increased abduction (i.e. decreased adduction as shown in Figure 9) [2].
@@
The effects of fatigue are also evident in waving movement, where the elbow of the right (i.e. waving hand) is rendered more extended, resulting in a limited range of motion (Section C in Figure 9 and Figure 8). Overall, the greater the decrease in percent of maximum exerted torques (i.e. greater $RC(t)$ as arises from Equation (8)) the more the influence of fatigue in the produced motions. [^4] Please refer to the videos in the supplemental material for more qualitative results regarding our framework.
Figure 8: Effects of fatigue at waving motion. The starting and midwave points of the movement are indicated by yellow markers, while the angle between the waving right hand and the starting point is denoted by red arrows.

### v2

@@ III-A Dataset & Data Processing
Fatigue-PINN leverages a PINN version of 3CC, namely 3CC-$\lambda$, for modeling the fatigue state of specific muscle groups/joint actuator torques (w.r.t. the animated motion) during open human movements. As shown in Fig. 1, our PINN 3CC-$\lambda$ is given as input the current active motor units $M_{A}$, which in our case are the joint torques of each DoF, $\hat{\mathbf{\tau}}_{t}^{N}$, produced by the ID model, and outputs the percentage of motor units in the complementary states, $M_{F}$ and $M_{R}$ in $MVC\%$. In contrast to previous work [44], we compute the $RC$ of each joint, based on the $M_{F}$ state multiplied by a $\lambda$ factor (Eq. (7)) that empirically describes the fatigue profile of each muscle/joint to better align joint fatigue in different movements and allow for joint-specific fatigue configurations. The $\lambda$ factor, ranging between 0 and 1, is strategically selected to mitigate the influence of fatigue on motion by a small percentage, hence producing smoother motions. This parameter is action-dependent and varies for each joint. Based on Eq. (8), joints whose fatigue profile is set to $\lambda=0$ will not contribute to the fatigued motion (fatigued torques will not be applied upon them), while a joint with $\lambda=1$, has its $RC$ reduced by the $M_{F}$ factor. In the latter case, the fatigued torques applied to the respective joints are completely dependent on the 3CC, i.e. the ones arising from the fatigue state. Subsequently, we apply $RC$ as a time-varying factor to simulate the gradual decaying of the maximum exerted joint torques to directly generate the fatigued ones, $\acute{\hat{\mathbf{\tau}}}_{t}^{N}$, as in [23, 44] (Eq. (8)):
@@
For instance, if $M_{F}=20\%$, at elbow joint at frame t, and $\lambda=0.6$ then from Eq. (7) we derive that the maximum exerted elbow torques peak at $RC(t)=88\%$ (instead of $RC(t)=80\%$), meaning that $\lambda$ factor, contributes in smoothly injecting joint-specific fatigue and producing more physically-plausible motions without intermittent (”step-like”) movements.
As depicted in Fig. 1, the 3CC-$\lambda$ model consists of five fully connected layers, being fed with the current time frame $t$ and current torque of each DoF, $\hat{\mathbf{\tau}}_{t}^{N}$ (i.e. $M_{A}$), and produces $M_{F}$ and $M_{R}$ of the same frame. Its loss, $L$ as formulated in Equations (9), (10), and (11), is comprised of two terms (in accordance with Fig. 1) i) the Mean Squared Error (MSE) loss of the neural network, $L_{NN}$, and ii) the physics-based loss, $L_{PB}$, based on Equations (3) and (1) that describe the transition between the active and fatigued state and the transition between the fatigued and the resting state, respectively:
@@ III-C Inverse/Forward Dynamics Model
The forward and backward sequence outputs are computed using Equations (12) and (13), which are the standard LSTM updating equations [66, 68, 67]. In particular, the output $\overrightarrow{h}_{t}$ of the forward sequence is computed iteratively in a positive time direction spanning from $t=T$ to time $t=1$. On the contrary, the backward loop is ”looking” at the past, thus, its output $\overleftarrow{h}_{t}$ is calculated in a negative time direction from $t=1$ to $t=T$ (see Fig. 3):
@@
where $x_{t}$, is the input data $x$ at time frame $t$, $U$ and $W$ are the weight matrices and $b$ is the bias vector. The input of the ID model is defined as $x_{t}=\mathbf{q}_{t}^{N}$, i.e. a vector of $N$ joint angles at frame $t$, while $x_{t}=\hat{\mathbf{\tau}}_{t}^{N}$ is the input for the FD model, i.e. a vector of $N$ fatigued torques at time $t$. The values of the weight matrices and the bias vector are updated (”learned”) during training.
Furthermore, the BiLSTM layer produces an output vector at each time frame $t$, $y_{t}$, which is computed using the calculated feedback sequences’ outputs, $\overrightarrow{h}_{t}$ and $\overleftarrow{h}_{t}$, as indicated by Eq. (14) [67].
@@
where in our case $g$ is a concatenating function used to combine the outputs of the two feedback sequences [68]. Given that $y_{t}$ is the output of the last BiLSTM layer in each of the ID/FD architecture, it is then passed to a fully connected layer (as illustrated in Figures 1 and 3). In our joint-specific framework, the fully connected layer in the ID model computes the current torque for one joint denoted as $\hat{\mathbf{\tau}}_{t}$, while the fully connected layer of the FD architecture produces a single fatigued joint angle, $\acute{\hat{\mathbf{q}}}_{t}$, at frame $t$.
As shown in Figures 1 and 3, the ID and FD models utilize a symmetric BiLSTM architecture consisting of five BiLSTM layers each one of them containing 128 units with linear activation on the first BiLSTM layer and ReLU activations on the rest hidden layers. Additionally, a Fully Connected layer with linear activation is included as the output of the ID/FD models. Similarly to the training of 3CC-$\lambda$, both ID and FD BiLSTM networks were trained in Python Keras, utilizing the Adam optimizer with a learning rate of $0.001$. The learning process was conducted with a batch size of 32 and an MSE loss function, over 1000 epochs, with early stopping enabled to prevent the model from over-fitting. Our joint-specific approach, meaning training a Fatigue-PINN model separately for each joint angle, may entail over-fitting. Producing predictions (either torques or angles) only for one joint requires extracting latent-space dependencies for one output, which results in faster convergence, i.e. in fewer epochs, rendering early stopping essential. Early stopping also tackles over-fitting issues due to the lack of motion capture data from multiple subjects.
The final models and hyperparameters were selected after assessing various architectures for our ID/FD BiLSTM models in terms of the number of BiLSTM layers and units, batch size, and learning rate. In Table I, the MSE values of these configurations are reported indicatively for the punching motion type, to assess the testing accuracy of every architecture. MSE measures the average squared difference between the ground-truth values and the predicted values. During the computation of this metric, the errors are squared, meaning that larger errors can significantly increase the MSE and indicate a deterioration in performance.
@@
TABLE I: Comparison Across ID/FD BiLSTM Architectures
Variations in Number of BiLSTM Layers (128 units, 32 batch size, lr = $0.001$)
Model 3 layers 4 layers 5 layers 6 layers
ID (MSE) $0.043$ $0.060$ 0.028 $0.11$
FD (MSE) $0.88$ $0.20$ 0.152 $0.27$
Variations in Number of Units (5 layers, 32 batch size, lr = $0.001$)
Model 64 units 128 units 256 units 512 unit
ID (MSE) $0.048$ 0.028 $0.053$ $0.036$
FD (MSE) $0.91$ 0.152 $0.21$ $0.26$
Variations in Batch Size (5 layers, 128 units, lr = $0.001$)
Model 16 batch size 32 batch size 64 batch size 128 batch size
ID (MSE) $0.047$ 0.028 $0.060$ $0.035$
FD (MSE) $0.65$ 0.152 $0.87$ $0.23$
Variations in Learning Rate (5 layers, 128 units, 32 batch size)
Model lr = $0.01$ lr = $0.001$ lr = $0.0001$ lr = $0.00001$
ID (MSE) $0.86$ 0.028 $0.032$ $0.050$
FD (MSE) $7.3$ 0.152 $0.18$ $0.21$
  • a
    A comparison between different ID/FD BiLSTM architecture and hyperparameter combinations in terms of Mean Squared Error (MSE). Evaluations on variations of a hyperparameter were conducted independently of the rest of the hyperparameters of an architecture. For example, when imposing changes in the number of BiLSTM layers on a model, the rest of the hyperparameters remain unchanged (e.g., 128 units, 32 batch size, lr = 0.001). All architectures were evaluated indicatively on the punching motion type. The architectures indicating the best performance are marked in bold. The abbreviation ”lr” stands for learning rate. All values are multiplied by $10^{-2}$.
From Table I it is clear that the architecture chosen for our ID/FD models, namely 5 BiLSTM layers with 128 units each, batch size = $32$, and learning rate = $0.001$, has the best MSE values compared to the rest of the models that we evaluated. It is worth mentioning that these MSE values are measured over the normalized data, ranging from 0 to 1. Therefore, an MSE of $0.00152$ resulting in an RMSE = $\sqrt{0.00152}\sim=0.0389$, shows that the predictions of our model differ on average by $0.0389$ from the ground-truth values. Thus, smaller MSE values indicate a minimal error and, as a result, a high accuracy. Overall, a small enough batch size enables the model to learn over small parts of the training dataset at each iteration, while a learning rate of $0.001$, which is commonly used in literature, offers a good frequency of updating the model’s weights without significantly increasing training time (smaller learning rate values result in updating weights more often, hence larger training times). Moreover, utilizing 128 units provides a good balance between accuracy and the number of training parameters as well as training time (i.e. the more units used for each layer, the more the parameters). Similar MSE values were yielded by testing the same architectures on throwing and waving motion types.
@@ IV Results
Figure 4: The impact of different levels of fatigue on punching (Frame 290, i.e. the moment of hit), throwing (Frame 240, at throw), and waving (Frame 100) motions. The levels of fatigue are defined w.r.t Residual Capacity as arises from Eq. (8). For instance, $30\%$ fatigue $\rightarrow RC(t)=70\%$, etc.
@@ IV-A Fatigue Effects in Animation
The proposed Fatigue-PINN framework was trained and tested upon open-type movements (as defined in this work), i.e. waving, punching, throwing, to assess its capability of effectively producing fatigued motion in various action classes (see Fig. 4). It is worth mentioning that we train a Fatigue-PINN framework for each one of the aforementioned motion types. Therefore, for each action, we performed cross-trial validation to enhance the robustness and accuracy of the model across different motion trials (i.e. emotional variations of the same movement).
We opted to apply fatigued torques on the upper body (shoulders, elbows, wrists, and lumbar), whose motion is crucial for perceiving the effects of fatigue in these particular open motions. Therefore, the fatigue profiles -$\lambda$ values- for the upper body joints are set between $0.6$ to $1$ across different motion types (punching, waving, and throwing), to slightly reduce the influence of fatigue (i.e. produce smoother movements), yet retain the effects of the fatigue state produced by the 3CC on the resulting motion. The $\lambda$ factor values for the lower body joints in all scenarios presented in this work are set to 0.
Our results are simulated in Unity Game Engine Environment [69], utilizing the Anime Girl Character [^4] Available online at: https://assetstore.unity.com/packages/3d/characters/humanoids/casual-1-anime-girl-characters-185076#publisher avatar from Unity Asset Store.
@@
In both motion frames depicted in Fig. 5, the character on the left exhibits fatigued movement ($RC(t)=\sim 55\%$) in all joints of both arms, while the right is the non-fatigued one. In punching motion (Fig. 5), the hand positioning is different than in the non-fatigued motion, rendering the fatigued character ”unguarded” to hits (see frame $100$ in Fig. 5). Moreover, in frame $290$ as shown in the right section of Fig. 5, another discrepancy in motion due to fatigue is apparent, with both arms performing a wider movement while hitting, thus, missing the target. These two effects on the kinesiology of the 3D character, guard drop, and non-accurate punches, are behaviors observed alongside the increase of perceived fatigue in boxing athletes [18, 19], while guard drops pose as a tactical strategy to mitigate fatigue by providing a small period of rest [18]. The latter, known as a pacing strategy, involves the adaptation of exercise intensity (decrease/increase effort accordingly) to minimize the impact of fatigue and thereby attain the desired outcome [17] (e.g. determine the winner in a boxing bout). Moreover, the ”unguarded” stance is a result of increased shoulder abduction, therefore, decreased shoulder adduction as seen in Section A of Fig. 9, and increased elbow extension (i.e. less flexed elbow as in Fig. 9). Notably, while shoulder adduction generally decreases, it increases at the moment of hitting impact (around frame $250$), resulting in a corresponding decrease in shoulder abduction. These findings align with those in [3]. Lumbar bending and rotation are reduced, with the character inclining to the left during striking with the right hand, while the upper body exhibits greater rotation to the right. These discrepancies in joint angles are reflected in the respective torques, which are expected to be reduced in the presence of fatigue by the 3CC-$\lambda$ model.
@@
Similar results are observed in throwing (see Fig. 6), with the right hand, the one performing the motion, being directed in a lower position than the one intended. The externally perceived impact of fatigue in this movement can be associated with the one in pointing tasks [70, 2], where one hand extends towards a specific position while the other remains relatively inactive. However, throwing is inherently more physically demanding, thereby we intuitively expect it to lead to a more pronounced fatigue effect. Regarding the changes in position as depicted in the right section of Fig. 6 (frame $240$), the shoulder joint positions of both arms are moving towards the non-throwing side, while the right shoulder is positioned higher at throw, which is in accordance with the findings in [70], and may provide compensation mechanisms for the increased shoulder elevation and elbow extension [70, 71]. Moreover, following fatigue, the shoulder and elbow joints of the right arm (the one performing the throwing/pointing task) adopt a more posterior position, indicating a posture that reduces shoulder torques [70] (Fig. 7), consistent with the predictions of the 3CC-$\lambda$ model.
Similar to the findings in [70, 2], the fatigued character in the left of Fig. 7 exhibits reduced shoulder flexion of the right hand at the moment of throwing (frame $240$). Akin to the results in [70], upper-body inclination towards the non-reaching arm’s side due to decreased lumbar bending angle (i.e. more flexion in the lateral plane towards the left side) is also evident at throw (Frame $240$ in Fig. 6). Moreover, in accordance with [2], greater lumbar rotation (i.e. increased lumbar left rotation angle) throughout throwing (frames 210300210-300) is apparent in fatigued movement as illustrated in Figures 6 and 9 (Section B). The above kinematic effects are attributed to elbow and/or shoulder fatigue, as observed in [70, 1, 2].
Furthermore, the right-hand elbow is less flexed (reduced elbow flexion angle as seen in Figures 6 and 9), which is considered a mean of compensation for both elbow and trunk fatigue as mentioned in [2], whereas elbow fatigue also influences shoulder abduction/adduction, leading to increased abduction (i.e. decreased adduction as shown in Fig. 9) [2].
@@
The effects of fatigue are also evident in waving movement, where the elbow of the right (i.e. waving hand) is rendered more extended, resulting in a limited range of motion (Section C in Fig. 9 and Fig. 8). Overall, the greater the decrease in the percent of maximum exerted torques (i.e. greater $RC(t)$ as arises from Eq. (8)) the more the influence of fatigue in the produced motions. [^5] Please refer to the video in the supplemental material for more qualitative results regarding our framework.
Figure 8: Effects of fatigue at waving motion. The starting and mid wave points of the movement are indicated by yellow markers, while red arrows denote the angle between the waving right hand and the starting point.